[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Mar 30, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Mar 30 18:42:43 UTC 2017


HSU Shift Log for Thu Mar 30 2017 14:42:36 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: GFS data is still very delayed.

Other Techincal issues: The Hurricane Hotline has been repaired, or at  
least it is no longer giving off that horrible noise.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170330/e3f71f27/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list